# Robust Policy Switching for Antifragile Reinforcement Learning for UAV Deconfliction in Adversarial Environments

### v1

@@ V-C Antifragile Behaivour
A RL system is said to exhibit antifragility under adversarial observation perturbations if, in response to increased perturbation intensity or out-of-distribution (OOD) attacks, the system adaptively improves its expected performance (reward) over time by leveraging the variability to refine its policy selection mechanism.
@@
Let $\mathcal{P}=\{\pi_{\alpha_{1}},\dots,\pi_{\alpha_{K}}\}$ be a finite set of robust policies trained under adversarial intensities $\alpha_{k}\in[0,1]$. Let $\kappa_{t}$ denote the policy selected at time $t$ by the DTS-based switching mechanism under an adversarial perturbation $\epsilon_{t}\notin\{\alpha_{1},\dots,\alpha_{K}\}$, i.e., an unseen attack. We consider the value shift satisfies the Lipschitz robustness assumption as in (1), with the reward proxy defined in (27) and DTS maintains a discounted Beta posterior as in (22). Then, under slowly evolving or piecewise-constant perturbation intensity $\epsilon_{t}$, the expected reward of the selected policy improves over time:
@@
$\lim_{t\to\infty}\mathbb{P}\left(\kappa_{t}=\arg\min_{k}d_{k}^{(t)}\right)=1.$ (34)
Let $k^{*}:=\arg\min_{k}d_{k}^{(t)}$ denote the optimal policy under the current (unseen) perturbation. Since DTS samples utilizes the optimal policy $k^{*}$, and $r_{k}^{(t)}=d_{\min}^{(t)}/d_{k}^{(t)}$, we have:
$\lim_{t\to\infty}r_{\kappa_{t}}^{(t)}\to 1\quad\Rightarrow\quad\lim_{t\to\infty}d_{\kappa_{t}}^{(t)}\to d_{\min}^{(t)}.$ (35)
By the Lipschitz assumption:
$\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]\geq\mathbb{E}[R(\pi_{0};\Phi_{t})]-L\cdot d_{\kappa_{t}}^{(t)}.$ (36)
Differentiating both sides w.r.t. $t$, we get:
$\frac{d}{dt}\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]\geq-L\cdot\frac{d}{dt}d_{\kappa_{t}}^{(t)}.$ (37)
$\frac{d}{dt}d_{\kappa_{t}}^{(t)}<0\quad\Rightarrow\quad\frac{d}{dt}\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]>0.$ (38)
@@ VI-A Numerical Implementation
The software implementation of the DDPG algorithm, used for the deconfliction of UAV navigation in the environment consisting of dynamic 3D obstacles, has been considered as per [IFDSCA]. During the training phase, the initial position of the UAV was randomly considered with a mean around $\left(0,2,5\right)$ and a variance between 0 and 1, while the target destination was fixed at $\left(10,10,5.5\right)$. Kinematic constraints, including a maximum ascent angle of $95\pi/$ 9 and a maximum descent angle of $-15\pi/36$, were applied when updating the UAV position. In addition, a protection radius of 1.5 units was enforced around the UAV. The dynamic obstacle used in training was restricted to follow a predefined path along the x-y plane, with its trajectory given by the following equation:
$\begin{gathered}\mathbf{x}_{\textup{obs}}\left(t\right)=\mathbf{x}_{\textup{obs}}\left(t-1\right)+2\cos\left(t\right),\\ \mathbf{y}_{\textup{obs}}\left(t\right)=\mathbf{y}_{\textup{obs}}\left(t-1\right)+2\sin\left(t\right).\\ \end{gathered}$ (39)
The reward function calculates the conflict penalty by augmenting the UAV’s radius with an additional buffer of 0.4 units. The vanilla agent and the adversarial action robust models in the ensemble were trained considering a single dynamic obstacle. The action robust RL model was trained according to the Algorithm 1, using the parameters specified in Table I.
@@
VI-B Training Results
Figure 5: Convergence plot for vanilla DDPG and adversarial action robust DDPG for various $\alpha$.
The training results, based on the cumulative reward for the episode, are illustrated in Figure 5. The action robust RL model for adversarial actions was trained using varying proportions of adversarial influence in the action space, denoted as $\alpha=\left\{0.1,0.2,0.3,0.4\right\}$. As shown in Figure 5, increasing $\alpha$ leads to greater uncertainty in the converged rewards, as reflected in the widening uncertainty bands. This trend is further confirmed by the mean entropy difference $\Delta\mathbf{H}$ in Figure 6, where the gap between exploration and exploitation rewards decreases linearly with increasing $\alpha$. To defend against evasion attacks during testing, it is crucial to train models with a less convex loss landscape. However, as the landscape of loss flattens, greater randomness is introduced into the policy space, contributing to increased uncertainty in the final rewards.
VI-C Distribution Shift against Adversarial Attacks
Figure 7: The distribution shift for the critic values for the action robust RL with $\alpha=\{0.1,0.2,0.3,0.4\}$ under normal operations and Frank Wolfe adversarial attack with $\epsilon=2.5.$
A detailed distributional analysis reveals that the shift in the state-action value distribution, i.e. $\mathbf{d}^{\alpha}_{\epsilon}$, exhibits a linear increase with increasing $\epsilon$ in all models. In particular, for $\alpha=0.3$, the magnitude of the distributional shift remains lower compared to other robust models. This trend is evident in Figure 7, where the mode of distribution of $\alpha=0.3$ is closely aligned with that of the standard DDPG model, in contrast to other robust policies.
Next, we examine the behavior of $\mathbf{p}_{\textup{true}}$ that normalizes the distribution shift $\mathbf{d}_{\epsilon}$ according to (20). The multiplicative factor $k_{\textup{mult}}$ that ensures a probability of 10% of selecting non-optimal models during deployment. A distinct variation in $p_{\textup{true}}$ is observed between low and high values of $\epsilon$. Specifically, for $\alpha=0.3$, $p_{\textup{true}}$ is the highest compared to DDPG and other robust action models. However, at $\epsilon=1.0$, $p_{\textup{true}}$ is maximized for $\alpha=0.2$, reinforcing the rationale behind employing DTS sampler instead of standard TS. A sharp transition in Bernoulli rewards is also expected between $\epsilon=0.5$ and $\epsilon=1.0$. Furthermore, as $\epsilon$ increases from 1.5 to 2.5, the relative decline in $p_{\textup{true}}$ amongst the action robust models diminishes, with $\alpha=0.3$ consistently yielding the highest value of $p_{\textup{true}}$. These findings suggest that the DTS should be trained to switch amongst the action robust models at a higher frequency to mitigate the impact of adversarially induced distributional shifts.
Figure 8: a) The Wasserstein distance comparison between the critic values of vanilla DDPG and various action-robust DDPG formulations under incremental adversarial attacks. (b) The distributional shift induced by adversarial attacks, represented through the parameters of the Bernoulli distribution used for sampling.
Figure 9: The number of conflict-free paths generated by the antifragile algorithm over 100 episodes while comparing different sampling strategies. PGD adversarial attacks were applied at (a) $\epsilon=2.0$ (b) $\epsilon=2.5$.
VI-D Results against PGD Adversarial Attack
We further examine navigation performance under adversarial attacks, specifically considering a Projected Gradient Descent (PGD) attack with $N_{\textup{steps}}=50$. In practical scenarios, such attacks require knowledge of the policy network governing the navigation of the UAV. This policy can be inferred using generalized adversarial imitation learning [ho2016generative] by exploiting unencrypted ADS-B data exchanges, which contain position and velocity information for both the UAV and dynamic obstacles. Using the extracted policy, an adversary can craft stealthy perturbations of a given magnitude $\epsilon$. The resulting test rewards, depicted in Figure 10, correspond to the initial randomized conditions between episodes.
The antifragile algorithm that uses the DTS sampler demonstrates superior performance, achieving higher testing rewards compared to the baseline algorithms. Furthermore, the variance in rewards is lower for the DTS-based approach, indicating increased stability. Safety performance is assessed by analyzing both the total number of conflicts and the proportion of conflict-free trajectories in the testing environment in Figures 11 and 12. The DTS algorithm significantly outperforms the robust and vanilla standard RL models in minimizing conflicts. However, while DTS marginally surpasses adversarial learning methods in overall performance, the latter exhibits a lower variance in conflict occurrences. Despite this, Figure 10 reveals that adversarial learning achieves lower overall rewards than the DTS-based method. This suggests that while adversarial learning improves collision safety, it does so at the expense of efficient path planning, thereby increasing vulnerability to alternative forms of attack.
Figure 10: The rewards obtained over 100 testing episodes in a multi-obstacle environment when evaluating the DTS-based antifragile algorithm against PGD adversarial attacks, in comparison to other robust RL techniques and vanilla DDPG.
Figure 11: The total number of conflicts encountered during UAV navigation under PGD adversarial attacks in a multi-obstacle environment, comparing the DTS-based antifragile algorithm with other robust RL techniques and vanilla DDPG over 100 episodes.
Figure 12: The total number of conflict-free paths generated during UAV navigation under PGD adversarial attacks in a multi-obstacle environment, comparing the DTS-based antifragile algorithm with other robust RL techniques and vanilla DDPG over 100 episodes.
VI-E Results against Spoofing Attacks for Hijacking
A separate analysis considers a different adversarial strategy, in which the position of the UAV is spoofed to facilitate hijacking [giray2013anatomy]. This attack is commonly used to manipulate position inputs, forcing UAVs to elongate flight paths through specific airspace zones, increasing exposure to hijacking attempts. To simulate this, the position of the UAV in each time step is perturbed with a uniform random bias $\mathcal{U}\left(0.04,0.06\right)$. The objective is to determine whether the UAV deviates significantly from its intended trajectory. This scenario exemplifies antifragility, in which the algorithm is explicitly trained against adversarial disturbances to adapt to previously unseen attack patterns. As illustrated in Figure 13, the DTS-based antifragile algorithm maintains an optimal trajectory with minimal deviation, unlike alternative methods that exhibit excessive compensation in response to adversarial perturbations. Specifically, the rewards for NR-MDP and PR-MDP are notably lower due to their reliance on fixed countermeasures. In contrast, the adaptive nature of the DTS sampler, trained in non-stationary Bernoulli rewards, enables effective path planning while mitigating deviations induced by spoofing attacks. Consequently, the proposed approach ensures improved resilience against hijacking attempts while preserving efficient navigation.
Figure 13: The rewards obtained over 100 testing episodes when evaluating the antifragile algorithm with a DTS sampler against other robust RL techniques and vanilla DDPG under spoofing attacks.
VII Conclusions
In this study, an antifragile reinforcement learning (RL) framework has been developed, employing a mechanism for switching between robust policies to enhance adaptability in the presence of adversarial observation attacks. DTS is utilized to optimally transition between robust models, mitigating the overall distributional shift in value estimation under adversarial perturbations. The theoretical framework is derived which states the optimality of DTS based switching and inducing antifragile behaivour against unseen adversarial attacks.
The proposed DTS-based antifragile RL approach is evaluated against progressively stronger adversarial attacks, including hijacking-induced spoofing, and benchmarked against conventional robust and standard RL techniques. The key findings of this study are as follows.
  • Non-stationarity of action robust policies under adversarial attacks: Analyzing various robust models under escalating adversarial attacks reveals that a single robust model does not maintain optimal performance across different attack intensities. This underscores the necessity of dynamic policy switching to achieve optimal resilience.
  • Enhanced Robustness Against Strong PGD Adversarial Attacks: The DTS-based antifragile RL algorithm outperforms standard and benchmark robust RL methods in countering stronger Projected Gradient Descent (PGD) attacks, particularly in terms of minimizing navigation path lengths with average reward improvement of 50% and increased average number of conflict-free trajectories by around 5 % with respect to adversarial RL and 10 % with respect to probabilistic robust RL.
  • Navigation Resilience to Spoofing Attacks: The proposed antifragile RL approach demonstrates superior resilience to previously unseen spoofing attacks compared to robust and standard RL models. This highlights the core principle of antifragility, in which a system trained under weaker adversarial conditions exhibits the capacity to counter unforeseen attack scenarios effectively.
Future research will focus on extending this antifragility framework to multi-UAV navigation scenarios, further enhancing its applicability in complex multi-agent adversarial airspace environments.
[…]
Appendix A Interfered Flow Field
Here, $d\left(\mathbf{P},O_{n}\right)$ represents the distance between the end of the path and the surface of the $N$-th obstacle, $\varrho_{0}$,$\varsigma_{0}$ represents the response coefficient of the obstacle. The dynamic obstacle speed threat is represented as
Appendix B Sampling Based Solution for Robust RL

### v3

@@ V-C Antifragile Behaviour
A RL system is said to exhibit antifragility under self-induced adversarial observation perturbations if, in response to increased perturbation intensity or out-of-distribution (OOD) attacks, the system adaptively improves its expected performance (reward) over time by leveraging the variability to refine its policy selection mechanism.
@@
Let $\mathcal{P}=\{\pi_{\alpha_{1}},\dots,\pi_{\alpha_{K}}\}$ be a finite set of robust policies trained under adversarial intensities $\alpha_{k}\in[0,1]$. Let $\kappa_{t}$ denote the policy selected at time $t$ by the DTS-based switching mechanism under an adversarial perturbation $\epsilon_{t}\notin\{\alpha_{1},\dots,\alpha_{K}\}$, i.e., an unseen attack. We consider the value shift satisfies the Lipschitz robustness assumption as in (25), with the reward proxy defined in (26) and DTS maintains a discounted Beta posterior as in (24). Then, under slowly evolving or piecewise-constant self-induced perturbation intensity $\epsilon_{t}$, the expected reward of the selected policy improves over time:
@@
The proof is provided in Appendix G. ∎
@@ VI-A Numerical Implementation
[…]
VI-B Training Results
Figure 4: Convergence plot for vanilla DDPG and adversarial action robust DDPG for $\alpha=\{0.1,0.2,0.3,0.4,0.5\}$.
The training results, based on the cumulative reward for the episode, are illustrated in Figure 4. The action robust RL model for adversarial actions was trained using varying proportions of adversarial influence in the action space, denoted as $\alpha=\left\{0.1,0.2,0.3,0.4,0.5\right\}$. As shown in Figure 4, increasing $\alpha$ leads to greater uncertainty in the converged rewards, as reflected in the widening uncertainty bands. This trend is further confirmed by the mean entropy difference $\Delta\mathbf{H}$ in Figure 5, where the gap between exploration and exploitation rewards decreases linearly with increasing $\alpha$. This is due to flattening of the loss landscape, due to greater randomness introduced in the policy space by the adversarial network, which in turn increased the uncertainty in the final rewards. For $\alpha=0.5$, there is no significant statistical difference between the rewards obtained during the exploration and exploitation phases. Consequently, robust policies corresponding to $\alpha=0.5$ are excluded from the set of policies considered for switching during the adaptation phase.
VI-C Distribution Shift against Adversarial Attacks
Figure 6: The distribution shift for the critic values for the action robust RL with $\alpha=\{0.1,0.2,0.3,0.4\}$ under normal operations and self-induced Frank Wolfe adversarial attack with $\epsilon=2.5.$
While a strict Lipschitz condition may not hold globally across all adversarial inputs, our empirical results provide evidence that small Wasserstein shifts in value distribution are correlated with the bounded changes in observed returns. In Figure 7b , we observe a monotonic, approximately linear relationship between the adversarial magnitude $\epsilon$ and the value distribution shift across the policies in the ensemble $\mathbf{d}^{\alpha}_{\epsilon}$ which suggests near-Lipschitz behaviour within the tested $\epsilon\in\left[0.5,2.5\right]$. In particular, for $\alpha=0.3$, the magnitude of the distributional shift remains lower compared to other robust models. This trend is evident in Figure 6, where the mode of distribution of $\alpha=0.3$ is closely aligned with that of the standard DDPG model, in contrast to other robust policies.
Next, we examine the behavior of $\mathbf{p}_{\textup{true}}$ that normalizes the distribution shift $\mathbf{d}_{\epsilon}$. The multiplicative factor $k_{\textup{mult}}$ that ensures a probability of 10% of selecting non-optimal models during deployment. A distinct variation in $p_{\textup{true}}$ is observed between low and high values of $\epsilon$. Specifically, for $\alpha=0.3$, $p_{\textup{true}}$ is the highest compared to DDPG and other robust action models. However, at $\epsilon=1.0$, $p_{\textup{true}}$ is maximized for $\alpha=0.2$, reinforcing the rationale behind employing DTS sampler instead of standard TS. A sharp transition in Bernoulli rewards is also expected between $\epsilon=0.5$ and $\epsilon=1.0$. Furthermore, as $\epsilon$ increases from 1.5 to 2.5, the relative decline in $p_{\textup{true}}$ amongst the action robust models diminishes, with $\alpha=0.3$ consistently yielding the highest value of $p_{\textup{true}}$. These findings suggest that the meta-policy based DTS is to be trained to switch amongst the action robust models at a higher frequency to mitigate the impact of adversarially induced distributional shifts.
Figure 7: a) The Wasserstein distance comparison between the critic values of vanilla DDPG and various action-robust DDPG formulations under incremental self-induced adversarial attacks. (b) The distributional shift induced by self-induced adversarial attacks, represented through the parameters of the Bernoulli distribution used for sampling.
Figure 8: The number of conflict-free paths generated by the meta-policy switching algorithm over 100 episodes while comparing different sampling strategies. PGD adversarial attacks were applied at (a) $\epsilon=2.0$ (b) $\epsilon=2.5$.
VI-D Results against PGD Adversarial Attack
We further examine navigation performance under adversarial attacks, specifically considering a Projected Gradient Descent (PGD) attack with $N_{\textup{steps}}=50$. The practicality of the attacks are described in Appendix B. The resulting test rewards, depicted in Figure 9, correspond to the initial randomized conditions between episodes.
The meta-policy switching algorithm demonstrates superior performance, achieving higher testing rewards compared to the baseline algorithms. Furthermore, the variance in rewards is lower for the DTS-based approach, indicating increased stability. Safety performance is assessed by analyzing both the total number of conflicts and the proportion of conflict-free trajectories in the testing environment in Figures 10 and 11. The DTS algorithm significantly outperforms the robust and vanilla standard RL models in minimizing conflicts. However, while DTS marginally surpasses adversarial learning methods in overall performance, the latter exhibits a lower variance in conflict occurrences. Despite this, Figure 9 reveals that adversarial learning achieves lower overall rewards than the DTS-based method. This suggests that while adversarial learning improves collision safety, it does so at the expense of efficient path planning, thereby increasing vulnerability to alternative forms of attack.
Figure 9: The rewards obtained over 100 testing episodes in a multi-obstacle environment when evaluating the DTS-based meta-policy switching algorithm against PGD adversarial attacks, in comparison to other robust RL techniques and vanilla DDPG.
Figure 10: The total number of conflicts encountered during UAV navigation under PGD adversarial attacks in a multi-obstacle environment, comparing the DTS-based meta-policy algorithm with other robust RL techniques and vanilla DDPG over 100 episodes.
Figure 11: The total number of conflict-free paths generated during UAV navigation under PGD adversarial attacks in a multi-obstacle environment, comparing the DTS-based meta-policy switching algorithm with other robust RL techniques and vanilla DDPG over 100 episodes.
VI-E Results against Spoofing Attacks for Hijacking
A separate analysis considers a different adversarial strategy, in which the position of the UAV is spoofed to facilitate hijacking as detailed in Appendix B. This attack is commonly used to manipulate position inputs, forcing UAVs to elongate flight paths through specific airspace zones, increasing exposure to hijacking attempts. To simulate this, the position of the UAV in each time step is perturbed with a uniform random bias $\mathcal{U}\left(0.04,0.06\right)$. The objective is to determine whether the UAV deviates significantly from its intended trajectory. As illustrated in Figure 12, the DTS-based meta-policy algorithm maintains an optimal trajectory with minimal deviation, unlike alternative methods that exhibit excessive compensation in response to adversarial perturbations. Specifically, the rewards for NR-MDP and PR-MDP are notably lower due to their reliance on fixed countermeasures. In contrast, the adaptive nature of the DTS sampler, trained in non-stationary Bernoulli rewards, enables effective path planning while mitigating deviations induced by spoofing attacks. Consequently, the proposed approach ensures improved resilience against hijacking attempts while preserving efficient navigation.
Figure 12: The rewards obtained over 100 testing episodes when evaluating the antifragile algorithm with a DTS sampler against other robust RL techniques and vanilla DDPG under high out-of-distribution attacks like GNSS spoofing.
Hence, we can conclude that the test rewards degrade smoothly with out-of-distribution adversarial attacks under non-switching baselines, while DTS mechanism preserves performance indicating that the degradation does not exhibit chaotic discontinuities in practice. The use of testing attacks as explained in SectionIV-D, is a security-stress test where our goal is to measure the adaptability of the system to different types of OOD perturbations. In real deployments, even partial adversarial knowledge can lead to policy appromixation attacks, control channel spoofing or physical-space decoy placement which is effectively abstracted from our synthetic adversarial observations.
VII Conclusions
In this study, a meta-policy adaptive framework has been developed, employing a mechanism for switching between robust policies to enhance adaptability in the presence of self-induced adversarial observation attacks. DTS is utilized to optimally transition between robust models, mitigating the overall distributional shift in value estimation under adversarial perturbations. The theoretical framework is derived which states the optimality of DTS based switching and inducing antifragile behaviour against unseen adversarial attacks.
The proposed DTS-based switching approach is evaluated against progressively stronger adversarial attacks, including hijacking-induced spoofing, and benchmarked against conventional robust and standard RL techniques. The key findings of this study are as follows.
  • Non-stationarity of action robust policies under adversarial attacks: Analyzing various robust models under escalating adversarial attacks reveals that a single robust model does not maintain optimal performance across different attack intensities. This underscores the necessity of dynamic policy switching to achieve optimal resilience. The ablation studies with respect to other bandit algorithms show that DTS maintains the best safety as compared to other algorithms.
  • Enhanced Robustness Against Strong PGD Adversarial Attacks: The DTS-based meta-policy switching framework outperforms standard and benchmark robust RL methods in countering stronger Projected Gradient Descent (PGD) attacks, particularly in terms of minimizing navigation path lengths with average reward improvement of 50% and increased average number of conflict-free trajectories by around 5% with respect to adversarial RL and 10% with respect to probabilistic robust RL.
  • Navigation Resilience to Spoofing Attacks: The meta-policy adaptive switching framework demonstrates superior resilience to previously unseen spoofing attacks compared to robust and standard RL models. This highlights the core principle of antifragility, in which a system trained under weaker adversarial conditions exhibits the capacity to counter unforeseen attack scenarios effectively.
In real-world scenarios involving highly dynamic or abruptly changing adversarial behaviors, these assumptions behind learning meta-policy switching may not hold. In such cases, additional mechanisms—such as online change-point detection [panda2025real] or reset-free meta-learning [lu2020reset] —should be integrated to maintain antifragility. Developing such extensions constitutes a key direction for future work.
Appendix A Interfered Flow Field
Here, $d\left(\mathbf{P},O_{n}\right)$ represents the distance between the end of the path and the surface of the $N$-th obstacle, $\varrho_{0}$$\varsigma_{0}$ represents the response coefficient of the obstacle. The dynamic obstacle speed threat is represented as
Appendix B Dual Threat Framework
We elaborate on the threat realism and relevance of each class below.
  • Observation-Space Attacks via Policy Inference (Grey-Box Threat Realism): These attacks emulate adversarial manipulation of the UAV’s state representation through inferred policy models, relying on the following feasible steps:
    • Step 1: In cooperative airspace environments, UAV telemetry—comprising position, velocity, and trajectory updates—is frequently broadcast over openly accessible communication channels such as ADS-B [manesh2017analysis, strohmeier2014security]. An adversary with passive eavesdropping capabilities can intercept these transmissions without detection.
    • Step 2: Leveraging the intercepted data, the adversary reconstructs the UAV’s input-output behavior using imitation learning frameworks such as Generative Adversarial Imitation Learning (GAIL) [ho2016generative] or behavior cloning [morales2004learning]. This yields a surrogate policy model $\hat{\pi}$ that approximates the UAV’s original navigation policy.
    • Step 3: Once $\hat{\pi}$ is obtained, the adversary can generate adversarial perturbations using gradient-based methods such as PGD [gupta2018cnn], effectively launching white-box style attacks on the approximated policy. These perturbations can be transmitted back to the UAV’s ADS-B IN interface, as malicious UTM operator, thereby influencing control decisions in a stealthy and coordinated manner.
    The feasibility of such inference-based attacks has been demonstrated in our prior work on adversarial mobility and surveillance [chu2024survey], where unencrypted trajectory data were exploited to construct targeted adversarial strategies. Since the induced distributional shifts due to PGD attacks resemble to those introduced by Frank-Wolfe-style perturbations, we categorize such attacks as low out-of-distribution (OOD) adversarial scenarios.
  • GNSS Spoofing Attacks (Black-Box Signal-Space Threats): In contrast to observation-space attacks, GNSS spoofing directly manipulates the UAV’s positional awareness by injecting falsified satellite signals. These attacks operate at the signal layer and do not require any access to the policy network or its gradients, thus constituting a fully black-box threat model. The attacker exploits pseudorange manipulation [panda2025real] to introduce consistent positional drift, misleading the UAV into computing paths based on erroneous geolocation data. Spoofing can be realized using low-cost software-defined radios (SDRs), constrained primarily by the proximity of the attacker to the UAV and the ability to overpower legitimate GNSS signals. This can lead to critical safety violations, including extended flight path, forced penetration into restricted zones, or increased collision risks. GNSS spoofing induces high-magnitude state-space shifts that are not gradient-driven and are thus semantically and statistically distinct from algorithmic perturbations from Frank-Wolfe adversarial attacks, we classify them as high OOD adversarial threats.
The inclusion of both adversarial observation perturbations and GNSS spoofing captures two fundamental axes of UAV system vulnerability: (1) control-layer adversaries, where the attack targets decision-making logic via state perturbations, and (2) sensor-layer adversaries, which compromise state estimation at the signal acquisition level. The proposed DTS-based policy switching framework enables robust adaptation across this threat spectrum, exhibiting resilience against both low and high OOD adversarial perturbations. In the following section, we present the theoretical guarantees underlying this adaptive robustness.
Appendix C Parameters for Training Robust Policies
@@
Appendix D Sampling Based Solution for Robust RL